• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Finn

    Log entry time 23:58:08 on August 4,2001

    Entry number 69267

    keyword=Shift summary

    Aug-04-2001 swing shift summary: Finn, Roche, Butuceanu
    ___________________________________________
    16:05: taking data on reference cell -9 Atm He pressure 118 psig
    16:27 at end of run 2567, pressure 103 psi
    16:30: request access to close He3 value on reference cell
    17:45 beam back- short run on BeO/C12 target with beam centered at (0,0)
    17:50 moving to polarized He3 cell start at 2uA
    18:00 first run with target cell shapiro
    18:30 raise beam current to 10uA-production running mode on pol He3 cell
    19:15 insert 6mstr collimators- continue production running
    22:15 complete lamba/2=out measurement
    22:15 begin ep measurement- ask for beam off for target movement
    22:30 complete target polatization measurements
    22:32 request beam lambda/2 plate change to IN position after consulting with halls B and C
    22:40 resume data taking with lambda/2 plate in IN position
    epr polarization is 48%


    _____________________________________________________
    Summary: poductive shift with stable beam with two interuptions for configuration changes

    6.3 hrs ABU
    1.7 hrs planned configuration changes